OpenStudy (anonymous):

The width of a rectangle is 6 feet more than its length. Find the width if the perimeter is 128 feet

OpenStudy (anonymous):

let the length be x 2(x+(x+6))=128 2x+6=64 2x=58 x=29 length=29 width=29+6=35